while playing game of billiards, your 0.50 kg cue ball travelling at 1.9 m/s glances off a stationary 0.40 kg board red billard balls moves at 1.3 m/s at an angle of 32 degrees clockwise from the cue ball s original path. what is the final speed of the cue ball? show all the work!


1
Expert's answer
2021-07-19T15:38:20-0400
"\theta=180-2(32)=116\degree\\v_1'=1.9\frac{\sqrt{0.5^2+0.4^2+2(0.5)(0.4)\cos{116}}}{0.5+0.4}=1.0\frac{m}{s}"